OpenStudy (anonymous):

a skier in the snow starts to descend from a height of 30m from rest, at a certain speed he gets to 20 m/s How high is he? Urgent please help!

OpenStudy (cj7529):

If you equate the loss in potential energy to the gain in potential energy you get mgh=0.5mv^2 m=mass g=acceleration due to gravity =9.82 ~10 h=height v = velocity divide both sides by mg h=[v^2]/[2g] and sub in your values

OpenStudy (shamim):

\[v ^{2}=u ^{2}+2g(30-x)\]

OpenStudy (shamim):

here v=final velocity=20m/s u=initial velocity=0m/s x=height of the skier frm the bottom of the snow when his speed is 20m/s

OpenStudy (shamim):

if u hv any confusion abt gravitational acceleration g then i will tell abt g more. plz ask me

OpenStudy (shamim):

\[or, 20^{2}=0^{2}+2g(30-x)\]

OpenStudy (shamim):

\[or, 400=2 \times9.8(30-x)\]

OpenStudy (shamim):

\[or, x=30-\frac{ 400 }{ 2\times9.8 }\]

OpenStudy (shamim):

\[or, x=9.6 m\]

OpenStudy (shamim):

so the height of skier frm the bottom of the snow is 9.6m

OpenStudy (shamim):

when the velocity of skier is 20 m/s